OpenStudy (aivantettet26):

Which of the following is the absolute maximum and minimum of the function: F(x) =2x^2 - 7x -10 on [-1, 3] respectively?

OpenStudy (dls):

Compute f'(x)

OpenStudy (aivantettet26):

this is the answer -1, -16 1/8

terenzreignz (terenzreignz):

Quadratic function... anyway, did you compute f'(x)?

OpenStudy (aivantettet26):

yes

OpenStudy (aivantettet26):

y' = 4x -7

terenzreignz (terenzreignz):

And when is this equal to zero?

OpenStudy (aivantettet26):

equate to zero?

terenzreignz (terenzreignz):

Yes

terenzreignz (terenzreignz):

So that means x = 7/4, yes?

OpenStudy (aivantettet26):

yes

terenzreignz (terenzreignz):

Does this fall within your interval?

OpenStudy (aivantettet26):

yes -1 and 3

terenzreignz (terenzreignz):

So make a note of that. Note that it's a minimum... why?

OpenStudy (aivantettet26):

coz it's positive

terenzreignz (terenzreignz):

What's positive? LOL Because its second derivative is positive at all points :P The second derivative is 4, right?

terenzreignz (terenzreignz):

Okay, so that point (x = 7/4) is POSSIBLY an absolute minimum. What's \[\Large f\left(\frac74\right)=\color{red}?\]

OpenStudy (aivantettet26):

replace all x with 7/4

terenzreignz (terenzreignz):

(Which, with a little brain-crunching, isn't really necessary) We know that in a quadratic function, the absolute minimum IS at the point where the derivative is zero, namely, at the vertex. Now check for its absolute maximum. Just check its value at the endpoints, -1 and 3.

terenzreignz (terenzreignz):

Are you still here?

OpenStudy (aivantettet26):

sorry bro.. My internet went wild